a rectangle with a perimeter of 26 inches has a length that is three inches longer than its width. What is the length and width of the rectangle?
Answer by ewatrrr(24785)

 
Hi,
.
2L + 2w = 26in
L + w = 13in L+=+w-3
(w-3) + w = 13
2w = 16
w = 8in and L = 5in 8-3
And...checking
L + w = 13in